QUESTION 4 a) A farmer wanted to find the relationship between the amount of fertilizer used and the yield of corn. He selected seven acres of his land on which he used different amounts of fertlizer to grow corn. The following table gives the amount (in kilograms) of fertilizer used and the yield (in bushels) of corn for each of the seven acres. Fertilizer used 120 80 100 70 Yield of corn 142 112 132 96 119 104 136 75 110 i) Calculate the correlation coefficient and describe the relationship between the measured variables. (6 marks) i) Plot the graph and derive the regression equation for the above data. (6 marks) ii) Predict the yield of corn for 105 kilograms of fertilizer applied. (2 marks) CzeTheit b) The Health Club poll asked adults whether they felt genetically modified food was safe to at. The result showed 35% felt it was safe, 52% felt it was not safe and 13% had no opinion. A random sample of 120 adults was asked the same question at a local county fair. A total of 40 people felt that genetically modified food was safe, 60 people felt it was not safe and 20 had no opinion. At the 0.01 level of significance, is there sufficient evidence to conclude that the proportions differ from those reported in the survey? (6 marks)
Answer #1

( a )

Let x = Fertilizer used , y = yield of corn

Find X⋅Y , X2 and Y2 as it was done in the table below.

X Y X⋅Y X⋅X Y⋅Y  
120 142 17040 14400 20164
80 112 8960 6400 12544
100 132 13200 10000 17424
70 96 6720 4900 9216
88 119 10472 7744 14161
75 104 7800 5625 10816
110 136 14960 12100 18496

Find the sum of every column to get: Σ X-643 , ΣΥ-841, ΣΧ.Y= 79152, Σ x? = 61169 , ΣΥ2-102821 Use the following formula to work out the correlation coefficient. 7-79152-643-841 r= 0.9814 7.61169 6432. 7-102821-8412

There is strong positive correlation between the measured variables

II)

150 140 130 +y=0:9027x +37.22 110 90 80 80 90 100 110 120 130 々.i) 60 70

( iii )

so regression equation from the above graph is

y = 0.9027x + 37.22

given x = 105

then

y = 0.9027 ( 105 ) + 37.22

y = 132.0035 ~ 132

y =132

( b )

Null hypothesis ( H0 ) : The proportion does not differ from those reported in the survey

Alternative hypothesis ( H1 ) : The proportion differ from those reported in the survey

The following table is obtained:

Categories Observed Expected (fo-fe)2/fe
Category 1 40 120*0.35=42 (40-42)2/42 = 0.095
Category 2 60 120*0.52=62.4 (60-62.4)2/62.4 = 0.092
Category 3 20 120*0.13=15.6 (20-15.6)2/15.6 = 1.241
Sum = 120 120 1.429

(2) Rejection Region Based on the information provided, the significance level is a 0.01, the number of degrees of freedom is df 3-1 2, so then the rejection region for this test is R- (x:x2 > 9.21) (3) Test Statistics The Chi-Squared statistic is computed as follows: Oi-E -= y. (01-b) = 0.095+ 0.092+ 1.241 = 1.429 (4) Decision about the null hypothesis Since it is observed that Χ-1.4296 χ-921, it is then concluded that the null hypothesis is not rejected

( 5 ) Conclusion : At alpha = 0.01 L.O.S there is no sufficient evidence to conclude that the proportion differ from those reported in the survey